OpenStudy (anonymous):

x for a graph that includes (0,6) and (1,15)

OpenStudy (anonymous):

if it's asking for an equation, the first point gives you the y intercept (when x=0, y=6) and the slope can be found by using the slope formula change in y/change in x slope = (15-6)/(1-0) = 9/1 = 9 then put it into slope intercept form y = slope*x + y-intercept y = 9x + 6